A water fountain on the ground sprinkles water all around it. If the speed of water coming out of the
fountain is v, the total area around the fountain that gets wet is

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is (pi*v^4)/(g^2)

Regards
Vikas TU
3 years ago
Rmax = (v2/g) sin2θ = v2/g
Area = πR2 = πv4/g2
….................................................................................................................